OpenStudy (anonymous):

Add 5 1 ____ + ____ 4z^3 10z^2

jimthompson5910 (jim_thompson5910):

\[\large \frac{5}{4z^3}+\frac{1}{10z^2}\] \[\large \frac{5*5}{5*4z^3}+\frac{1}{10z^2}\] \[\large \frac{25}{20z^3}+\frac{1}{10z^2}\] \[\large \frac{25}{20z^3}+\frac{2z*1}{2z*10z^2}\] \[\large \frac{25}{20z^3}+\frac{2z}{20z^3}\] \[\large \frac{25+2z}{20z^3}\] So \[\large \frac{5}{4z^3}+\frac{1}{10z^2}=\frac{25+2z}{20z^3}\]

OpenStudy (anonymous):